# Chain-structured neural architecture search for financial time series forecasting

### v1

@@ 5.2 Reinforcement learning approach
The reinforcement learning method consists of treating the neural architecture search problem as a reinforcement learning problem. We construct an iterative approach where the parameters (and structure) of the neural architecture are initially chosen randomly, just as in the case of Bayesian optimization. After the neural network is constructed, it is trained for a number of periods using a train and validation time-series dataset, sometimes even a number of times in order to correctly assess its performance in predicting the time series, and exclude issues relating to the choice of initial random seed. At the end of the run, a reward function is calculated, which represents the incremental increase in performance. In this way choices of parameters which produce “good” neural networks are rewarded, whereas those which produce badly-performing neural networks are penalized. We let the system evolve based for a number of iterations (reinforcement learning trials) and at the end we obtain the best overall neural network configuration.
The specific method we used introduced in [ZL17] uses an RNN as a controller, trained with REINFORCE. An open-source implementation of the controller RNN to use with image data is available at [Maj18], which we adpated to work with our time series data.
The reinforcement learning approach has several advantages compared to the Bayesian optimization: it is able to capture all the dynamics and interdependencies between the parameters, something that is difficult with classic bayesian optimization without testing every possible combination of parameters. Moreover we can more easily use customized reward functions (for example, in one case the reward function used was the average of the AUC and F1 scores) which can better capture the performance of the neural network in the presence of imbalanced datasets.
@@ 5.3 Hyperband
Hyperband is a method introduced in 2018 by Li et al., bringing in a clever tactic known as ‘bandit’ to make the hyperparameter optimization process very effective [${}^{+}$18]. It explores various combinations, paying more attention to the promising ones and discarding the under-performers.
@@
What makes Hyperband notable are its efficiency, scalability, and the ability to use multiple GPUs simultaneously for faster optimization. It is compatible with popular machine learning libraries like Scikit-learn and Keras, embedded in the KerasTuner framework [${}^{+}$19].
@@
Hyperband is designed for the efficient exploration and optimization of hyperparameter configurations. It selectively targets promising models in the search space, optimizing computational resources and time. The algorithm employs successive halving, a subprocess that involves eliminating less promising models and further training the survivors.
@@
In successive halving, you provide a list of models ($models$), the total epochs needed for full training ($max\_epochs$), the initial training epochs for all models ($start\_epochs$), and a factor determining the speed of reducing models and increasing training epochs for the survivors ($factor$). This process iterates until all models are fully trained, resulting in the identification of the best model and its validation loss.
@@
To achieve a balance between exploration and optimization, Hyperband executes successive halving multiple times, labeling each iteration as a bracket. The number of brackets is determined by a parameter, $s\_max$, approximately equal to $log(max\_epochs,factor)$. In each bracket, the number of models decreases by a factor, and new models are generated. Hyperband ensures a robust exploration by randomly generating models for each bracket, preventing too few models in a bracket. The initial training epochs for each bracket strictly increase by a factor.
@@
Practically implemented in KerasTuner as the Hyperband tuner, it is available for single or multiple GPUs for parallel tuning. Users can specify $factor$ and $max\_epochs$, along with the number of brackets using $hyperband\_iterations$. This allows control over the search time while efficiently exploring the hyperparameter space [SJH22].
@@ 6 Results
The data was split into (in historical order): first $70\%$ for training, the following $20\%$ for validation used in the evaluation during optimization runs, and the last $10\%$ for final testing on unseen data. When the optimization studies were completed, the architectures performing best on the validation dataset were selected and evaluated on the test set. This involved retraining and testing each network $50$ times, and average performance was compared. This was done for each search strategy, architecture type and dataset combination.
@@
The US dataset was most challenging for us. Although tuned 1d CNNs gave an AUC score of $0.6\pm 0.02$ on validation data (not used for training) in our repeated testing, and an individual lucky random seed gave an $AUC$ of $0.58$ on test data, on average no architecture could achieve $AUC$ over $0.5$ on the test dataset. We believe something very significant happened in the US market in the last year or so of our data that changed market dynamics.
@@ 6.1 Architectures and search strategies compared
Figure 1: Best performing architectures selected by each search strategy on each dataset. Every point represents average $AUC$ score and standard deviation on test data after retraining the selected architecture $50$ times. Type of the neural network chosen by the search strategy is displayed above each point, search strategies are color-coded.
For search strategies, overall the hyperband method and bayesian optimization showed better performance than the reinforcement learning based approach, but all three were very close to each other, well within standard deviation as shown in Figure 1. Optimization time was very similar between the three methods, mostly determined by the architecture type, total number of trials, repetitions per trial (to minimize variation due to different random seeds), and the number of epochs to train each neural network. These were set manually by us to 300, 15, and 80 respectively, for a total optimization time of about 12 hours per architecture type using a single Nvidia Quadro RTX 4000 GPU, with feedforward networks training faster than CNNs and LSTMs due to their lower complexity.
Practically, the reinforcement learning approach required the most work to adapt from the open-source implementation of the method in [ZL17], and has many parameters of its own to be optimized for best performance. In contrast, both the hyperband implementation in KerasTuner and the Bayesian optimization from Optuna are easy to use and well-supported. The Optuna implementation is flexible in that nearly the same code can be used to optimize hyperparameters of a different kind of solver other than neural networks, such as decision trees-based algorithms popular for time series prediction. Meanwhile, KerasTuner implementation of hyperband has advantage of easy parallelization over several GPUs (although we didn’t have the hardware to make use of it).
@@ 7 Discussion
Overall, we didn’t observe significant convergence over time in the hyperparameter optimization. An example history of optimization plot is shown in Figure 2. This was a general problem present for all neural architecture types, optimization strategies and datasets.
@@
We believe all search strategies were working as they should, but the main problem of financial time series prediction on our datasets was just too difficult to decisively improve upon through neural architecture search alone. Many steps were taken to improve overall performance. In the beginning the networks were heavily overfitting, where we struggled to get any metrics above $0.5$ even on the validation data. The feature reduction measures such as removing time-derived features and PCA discussed in Section 2 helped mitigate this. Unfortunately, this still did not always translate to good performance on the test dataset, even though we heavily used dropout layers for regularization. We also tried converting the main binary clasification problem to regression, but this did not result in substantial improvement.
@@
As a consequence, simply taking the final suggested architecture by either search strategy didn’t give the best results. The best performing ones were also not always the best choice due to the random seed variance discussed in Section 4. Even though each network configuration was trained and tested $15$ times during the optimization process, training and testing the best architectures more (e.g., 50) times sometimes showed worse performance, indicating that the issue of random seed variation was not fully resolved in our approach.
@@
Nevertheless, all search strategies were still successful in highlighting best choices for network parameters, at least for the validation set, which is most evident from slice plots as in Figure 3. For bayesian optimization, the best parameters for each architecture were read off slice plots like this. However we note that following this approach somewhat negates the main advantages in going for more advanced search strategies as compared to simple random search.
@@
Figure 3: Slice plot for bayesian optimization of LSTMs on the US dataset. Each point represents average $AUC$ score on the validation dataset over $15$ runs for the same network configuration. It is clear that setting chunk_length (i.e. the length in time of a sequence passed to the LSTM for individual prediction) to $10$ gives the best results, while other parameters are less impactful.
@@
To combat the big variance between networks starting from different random seeds (see Section 2 and Figure 1), we tried filtering the outputs, disregarding those with probabilities close to $0.5$ (interpreted as those where the network ‘was not sure’), a standard approach [GBC16]. Although this improved overall performance a little on average, it instead greatly increased the variance. For a fixed architecture, selecting networks trained from random seeds giving good performance on validation set did not translate to better performance on the test set. Our best suggestion to combat the variance is to use an ensemble model built from many incarnations of a chosen network architecture.
8 Future work
@@
Performing the same studies on more and publicly available time-series datasets, especially non-financial data would shed more light whether our findings were specific to our data or a more general feature.
@@
In addition, cell-based and hierarchical search spaces [${}^{+}$23] could be explored in the future in the context of time-series forecasting. These are based on searching over structural blocks within a network piece by piece instead of navigating through a grid-like space as in the standard approach we took. Finally, one-shot NAS techniques such as supernet-based methods (differentiable and not) and hypernetworks [${}^{+}$23] could be explored. With one-shot techniques, one trains a single (massive) super/hyper-network, which can be used to subsample and evaluate many smaller networks to find the optimal one, without re-training. This is opposed to the standard approach of training and testing many architecture configurations that we took.

### v2

@@ 4.2 Reinforcement learning approach
An alternative iterative approach consists of treating the neural architecture search problem as a reinforcement learning problem. Action of the agent at time step $t$ is selecting a specific neural network configuration. The reward is the performance of the selected model which leads to the next action. In this way choices of parameters which produce “good” neural networks are rewarded, whereas those which produce badly-performing neural networks are penalized. We let the system evolve for a number of iterations and at the end obtain the test results for all configurations explored.
The specific method we used introduced in [7] uses an RNN as a controller, trained with REINFORCE. The output sequence of the RNN controller is a string encoding the hyperparameters of a neural network to be trained and tested. The authors claimed that compared with Bayesian optimization, the reinforcement learning method is more general and more flexible [7]. The algorithm is presented in Algorithm 2. An open-source implementation of the controller RNN to use with image data is available at [22], which we adapted to work with our time series data.
Algorithm 2 Reinforcement learning NAS
0:  A search space $S$ of neural network hyperparameters, number of trials $n_{trials}$
0:  Results $R$ for each hyperparameter configuration tested
1:  Initialize the controller RNN $C$ with random weights $\theta_{C}$
2:  for $i=1$ to $n_{trials}$ do
3:     Evaluate $C$ to get an architecture configuration $s$ from $S$
4:     Train and test a neural net with configuration $s$, update results $R$
5:     Use the test metric as the reward signal, compute the policy gradient and update the controller weights $\theta_{C}$ of $C$
6:  end for
7:  return  $R$
@@ 4.3 Hyperband
The Hyperband introduced in 2018 by Li et al. [23], is a hyperparameter optimization method fundamentally different from both Bayesian optimization and the reinforcement learning approach. While the latter both use a clever technique to suggest the next neural network configuration to fully train and test, the Hyperband method selects many configurations randomly, instead cleverly allocating resources only to the ones looking promising early on in their training.
@@
What makes Hyperband notable are its simplicity, efficiency, scalability, and the ability to use multiple GPUs simultaneously for faster optimization. It is compatible with popular machine learning libraries like Scikit-learn and Keras, embedded in the KerasTuner framework [24].
@@
Hyperband is designed for the efficient exploration and optimization of hyperparameter configurations. Algorithm 3 employs successive halving, a subprocess that involves eliminating less promising models and further training the survivors.
@@
In successive halving, you provide the hyperparameter search space $S$, the total number of epochs $epochs_{\text{max}}$ needed for full training of a neural net, and a factor $\eta>1$ (usually $\eta=2$ or $3$ or $4$) determining the speed of reducing models and increasing training epochs for the survivors. This process iterates until the best performing models are fully trained, resulting in the identification of the best model and its validation loss.
@@
To achieve a balance between exploration and optimization, Hyperband executes successive halving multiple times, labeling each iteration as a bracket. The number of brackets is determined by a parameter $s_{\text{max}}=\left\lfloor\log_{\eta}(epochs_{\text{max}})\right\rfloor$. In each bracket, the number of models decreases by a factor, and new models are generated. Hyperband ensures a robust exploration by randomly generating models for each bracket, preventing too few models in a bracket. The initial training epochs for each bracket strictly increase by a factor.
Algorithm 3 Hyperband NAS
0:  Maximum number of epochs $epochs_{\text{max}}$, hyperparameter search space $S$, reduction factor $\eta>1$
0:  Set $R$ of best hyperparameters configurations maximizing the test metric
1:  Compute the maximum number of brackets: $s_{\text{max}}=\left\lfloor\log_{\eta}(epochs_{\text{max}})\right\rfloor$
2:  for  each bracket $s=s_{\text{max}},s_{\text{max}}-1,\ldots,0$ do
3:     Set the number of configurations $n_{configs}=\left\lceil\frac{s_{\text{max}}+1}{s+1}\cdot\eta^{s}\right\rceil$ in this bracket
4:     Set the initial number of epochs of training per configuration $n_{initial}=\frac{epochs_{\text{max}}}{\eta^{s}}$
5:     Randomly select $n_{configs}$ hyperparameter configurations from $S$ in a set $T$
6:     for  $i=0,1,\ldots,s$ do
7:        Train each configuration from $T$ for $n_{initial}*\eta^{i}$ epochs. Test and save the results.
8:        Select the top $\left\lfloor\frac{n_{configs}}{\eta^{i}}\right\rfloor$ configurations of $T$ based on the test metric, discard the rest from $T$
9:     end for
10:     Save the best hyperparameter configurations from $T$ in a set $R$
11:  end for
12:  return  the set $R$ of best hyperparameters
@@
Practically implemented in KerasTuner as the Hyperband tuner, it is available for single or multiple GPUs for parallel tuning. Users can specify $\eta$ and $epochs_{\text{max}}$. This allows control over the search time while efficiently exploring the hyperparameter search space [25].
A disadvantage of the Hyperband method is that it will quickly discard the architectures that performed badly early on in their training. Sometimes these architectures would outperform the others that did better early on if allowed to train until completion.
5 Methodology
5.1 Data preprocessing
To circumvent the high dimensionality of our data ($\sim$1000 features with under $4000$ samples), the number of features was reduced approximately by a factor of $3$ by removing features in the prepared datasets ‘time-derived’ from original features, such as by lagging original features by fixed time steps or taking the mean over the same time steps. Such time-derived features are standard in the industry [5], but are fundamentally used because the algorithms applied to the data can only take in one row of data as input. All the more sophisticated neural network architectures take in a sequence of rows of data, making time-derived features redundant. See Section 3 for more details. The datasets were then normalized before applying principal component analysis (PCA) to further reduce the number of features.
5.1.1 Principal component analysis
PCA is a dimensionality reduction technique widely used to transform high-dimensional data – especially when the features are highly correlated – into a lower-dimensional space by identifying the directions (principal components) along which the data varies most [1]. Let $X$ be our training data matrix of size $n_{samples}~{}\times~{}n_{features}$. PCA is an orthogonal linear transformation of the real inner product feature space that maps the data $X$ to a new coordinate system. In this transformed system, the direction of maximum variance by some scalar projection of the data aligns with the first coordinate (known as the first principal component), the direction of the second-highest variance aligns with the second coordinate, and so on. Moreover, the transformed columns of $X$ are linearly uncorrelated with each other.
It turns out that such transformation is given by a matrix $W$ whose columns are the eigenvectors of $X^{T}X$, ordered by their descending eigenvalues, which tell us how much of the variance in the dataset is described by the corresponding principal component [26]. By retaining only the first $k$ components with the highest variance, i.e., applying a transformation given by the first $k$ columns of $W$, we transform $X$ into a $n_{samples}~{}\times~{}k$ matrix, effectively capturing most of the information while reducing dimensionality, noise, and redundancy. A scree plot is a plot of the cumulative explained variance (i.e., cumulative sum of the eigenvalues, usually normalized so that the total sum is 1) against the component number. Looking at the scree plot, a good value of $k$ that balances dimensionality reduction and information retention can be identified (see Figure 1 for an example). We tried a couple values of $k$ for each dataset and architecture type before proceeding with the main NAS studies.
Figure 1: Scree plot for the Japan training dataset after time-derived features were removed. The curve saturates around $150$ components, suggesting that $k=150$ is a good choice to keep most of the information while reducing the dimensionality significantly.
5.2 Metrics
There are many evaluation metrics applicable to the binary classification problem. Although common for other classification tasks, accuracy ($ACC$) is not a good metric as a trivial predictor that always gives the same output no matter the input can have high accuracy in the case of an unbalanced dataset. Balanced accuracy ($bACC$), the average between the true negative and true positive rates, mitigates this problem. Another popular alternative is the $F_{1}$ score, the harmonic mean of precision and recall. The area under the receiver operating characteristic (ROC) curve, or $AUC$ for short, is another good choice. It is hard to say which one is better, and a good strategy could mix a combination of these. We mostly focused on $AUC$ as we found it to be the hardest metric to optimize, but we paid close attention to $bACC$ and $F1$ too and report on all three in Section 6.
5.3 Random seed variation mitigation
Neural networks, in particular in the context of financial time series forecasting suffer from comparatively high variance depending on the random seed chosen during training due to their stochastic nature [5, 6, 4]. Even identical neural network architectures, trained in exactly the same way but with different random seeds can have vastly different performance on the test dataset, e.g. an $AUC$ of $0.56$ on a ‘good’ random seed and $0.48$ (i.e., worse than random) on a ‘bad’ one. This is a problem for NAS as what seemed like a good (or bad) neural network configuration might have just been a lucky (or unlucky, respectively) seed. To mitigate this problem, we modified Algorithms 12, and 3 to train and test each architecture configuration chosen $15$ times with different random seeds, averaging the test metrics and using the averages as the result for the configuration chosen. At the end of optimization, the best performing configuration was trained and tested $50$ times (again, with different random seeds every time) to further eliminate random seed variation.
5.4 Experimental setup
We started with $7$ different broad architecture types: FFNN, depthwise 1D CNN, ordinary 1D CNN, 2D CNN, simple RNN, LSTM, and TFT, but for final optimization studies only FFNN, ordinary 1D CNN, and LSTM were used, as discussed in Section 3. Each of them has their own search space of hyperparameters determining the exact network architecture.
The data was split into (in historical order): first $70\%$ for training, the following $20\%$ for validation used in the evaluation during optimization runs, and the last $10\%$ for final testing on unseen data. Normalization and PCA transformations were fitted on the training datasets and applied to training, validation, and test data.
Each of the three search spaces was explored by each of the three search strategies described in Section 4 (modified by repeated training, testing, and averaging as discussed above) on each of the three (modified) datasets, resulting in $27$ studies total.
[…]
When the optimization studies were completed, the 27 architectures performing best on the validation datasets were selected and evaluated on the test set. For these final tests we retrained and tested each network 50 times and averaged the test metrics to minimize the variance coming from different random seeds.
@@ 6 Results
The US dataset was most challenging for us. Although tuned 1d CNNs gave an AUC score of $0.6\pm 0.02$ on validation data (not used for training) in our repeated testing, and an individual lucky random seed gave an $AUC$ of $0.58$ on test data, on average no architecture could achieve $AUC$ over $0.5$ on the test dataset. We believe something very significant happened in the US market in the last year or so of our data that changed the market dynamics completely.
@@ 6.1 Architectures and search strategies compared
Figure 2: Best performing architectures selected by each search strategy on each dataset. Every point represents average $AUC$ score and standard deviation on test data after retraining the selected architecture $50$ times. Type of the neural network chosen by the search strategy is displayed above each point; search strategies are color-coded.
For search strategy, overall the hyperband method and Bayesian optimization showed better performance than the reinforcement learning based approach, but all three were very close to each other, well within standard deviation as shown in Figure 2. Optimization time was very similar between the three methods, about 12 hours per architecture type using a single Nvidia Quadro RTX 4000 GPU, with feedforward networks training faster than CNNs and LSTMs due to their lower complexity.
Practically, the reinforcement learning approach required the most work to adapt from the open-source implementation of the method in [7], and has many parameters of its own to be optimized for best performance. In contrast, both the hyperband implementation in KerasTuner and the TPE Bayesian optimization from Optuna are easy to use and well-supported, providing many options for both optimization itself and performance monitoring. The Optuna implementation is flexible in that nearly the same code can be used to optimize hyperparameters of a different kind of solver other than neural networks, such as decision trees-based algorithms popular for time series prediction. Meanwhile, KerasTuner implementation of hyperband has advantage of easy parallelization over several GPUs (although we did not have the hardware to make use of it).
@@ 7 Discussion
Overall, we did not observe significant convergence over time in the hyperparameter optimization. An example history of optimization plot is shown in Figure 3. This was a general problem present for all neural architecture types, optimization strategies and datasets.
@@
We believe all search strategies were working as they should, but the main problem of financial time series prediction on our datasets was just too difficult to decisively improve upon through neural architecture search alone. Many steps were taken to improve overall performance. In the beginning the networks were heavily overfitting, where we struggled to get any metrics above $0.5$ even on the validation data. The feature reduction measures such as removing time-derived features and PCA discussed in Section 5.1 helped mitigate this. Unfortunately, this still did not always translate to good performance on the test dataset, even though we heavily used dropout layers for regularization. We also tried converting the main binary classification problem to regression, but this did not result in substantial improvement.
@@
As a consequence, simply taking the final suggested architecture by either search strategy did not give the best results. The best performing ones were also not always the best choice due to the random seed variance discussed in Section 5.3. Even though each network configuration was trained and tested $15$ times during the optimization process, training and testing the best architectures more (e.g., 50) times sometimes showed worse performance, indicating that the issue of random seed variation was not fully resolved in our approach.
@@
Nevertheless, all search strategies were still successful in highlighting best choices for network parameters, at least for the validation set, which is the most evident from slice plots as in Figure 4. For Bayesian optimization and the reinforcement learning approach, the best parameters for each architecture were read off slice plots like this.
@@
Figure 4: Slice plot for Bayesian optimization of LSTMs on the Japan dataset. Each point represents average $AUC$ score on the validation dataset over $15$ runs for the same network configuration. It is clear that setting chunk_length (i.e. the length in time of a sequence passed to the LSTM for individual prediction) to $10$ gives the best results, while the other parameters are less impactful.
@@
To combat the big variance between networks starting from different random seeds (see Section 5.1 and Figure 2), we tried filtering the outputs, disregarding those with probabilities close to $0.5$ (interpreted as those where the network ‘was not sure’), a standard approach [1]. Although this improved overall performance a little on average, it instead greatly increased the variance. For a fixed architecture, selecting networks trained from random seeds giving good performance on validation set did not translate to better performance on the test set. Our best suggestion to combat the variance is to use an ensemble model built from many incarnations of a chosen network architecture.
7.1 Future work
@@
Performing the same studies on more and publicly available time-series datasets, especially non-financial data would shed more light on whether our findings were specific to our data or a more general feature.
@@
In addition, cell-based and hierarchical search spaces [2] could be explored in the future in the context of time-series forecasting. These are based on searching over structural blocks within a network piece by piece instead of navigating through a grid-like space as in the standard approach we took. One-shot NAS techniques such as supernet-based methods (differentiable and not) and hypernetworks [2] could be explored. With one-shot techniques, one trains a single (massive) super/hyper-network, which can be used to subsample and evaluate many smaller networks to find the optimal one, without re-training. This is opposed to the standard approach of training and testing many architecture configurations that we took. Finally, genetic algorithms have proved to be very successful for NAS on image datasets [27] and should be explored in the context of time series too.
Statements and Declarations
  • Funding
    Funded by InnoSuisse (Project number: 54228.1 IP-ICT). First author was also supported by HEIG-VD projects HES-SO-340375, HES-SO-117951, and HES-SO-122845.
  • Competing interests
    The authors declare no competing interests.
  • Data availability statement
    The datasets presented in this article are private customer data provided by project partner Predictive Layer SA and are thus not readily available.
  • Author contributions
    Conceptualization, Stephan Robert-Nicoud, Biagio Nigro and Denis Levchenko; Methodology, Denis Levchenko; Software, Denis Levchenko, Biagio Nigro, Shabnam Ataee and Efstratios Rappos; Validation, Denis Levchenko, Efstratios Rappos and Shabnam Ataee; Investigation, Denis Levchenko, Biagio Nigro, Efstratios Rappos and Shabnam Ataee; Resources, Biagio Nigro and Stephan Robert-Nicoud; Writing – Original Draft Preparation, Denis Levchenko, Biagio Nigro, Shabnam Ataee, Efstratios Rappos; Writing – Review and Editing, Denis Levchenko, Efstratios Rappos; Supervision, Stephan Robert-Nicoud; Project Administration, Biagio Nigro and Stephan Robert-Nicoud; Funding Acquisition, Stephan Robert-Nicoud.